OpenStudy (anonymous):

The ordered pairs (1,16 )(2,25)(3,36)(4,49)(5,64) represent a function what Is a rule that represents this function

OpenStudy (anonymous):

A. y=x^2 B. y=x+15 C. y=2(x=3) D. y=(x=3)^2

OpenStudy (campbell_st):

well all the y values for the points are square numbers.... so that eliminates B and C so input x = 1 into A and D and see which givens y = 16 then to check use x = 3 to see which gives y = 36